UT Mar 03 2021




Night Partition

  • Night split with PEPSI/PETS - handover 23:30 to LBTO Eng LUCI-SOUL-AO
  • Moon rises 23:22 - through twilght (79%)
  • 1st half PEPSI/PETS
  • 2nd half LUCI-1 SOUL-AO

Issues:

  • MODS-2 is offline until further notice

Calibrations to be done

  • LUCI-1 & LUCI-2 G210 CWL 2.12 Arcs and Flats (Rothberg) for QSO Test
    • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/H-band_Library/

Night Plan Eng Time

AO Engineering Time - highest priority - (LBTO AO and Arcetri)
  • LUCI-SOUL AO tasks - test 6 different PAs per object
  • DONE LUCI-SOUL Arcetri tests
    • Seeing > 1.4" will not be useful for performance characterization
    • if seeing is steady above 1.4" above we can release the telescope for other ENG work
    • closed loop on AO ref stars (on-axis) to collect SR and telmetry
      • Complete collection of performance measurements
      • Alfio will be on-line to update settings at hand-over to impove HO1
    • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021//AO_Testing/AO_Testing_V22.xml
      • There are 33 stars with 6 different PAs (ignore Star 14) - will have to extract from OT before sending preset
    • one or more stars in the following categories
      • 14.5 < Rmag < 15.5 Fast Bin 3
      • 15.8 < Rmag < 16.5 Slow Bin 3
      • 16.51 < Rmag < 17.0 Fast Bin 4
      • 17.0 < Rmag < 18.0 Slow Bin 4
      • Rmag > 18.0 (as far as we can go....)
    • Procedure:
      • Preset with AOS in intervention mode and complete all the steps, including GOPT (enable NCPA for B3 only)
      • Double check that the DIMM is properly measuring seeing conditions
      • Check the flux on LUCI (Ks), and find the minimal DIT providing a good signal (max 20s in order to avoid PSF drifts inthe DIT)
      • Set the LUCI DIT in the Variable inspector of the AO system
      • Acquire the TN including the "PISCES PSF" with the required repetitions for a PSF that can provide a good measurement of the PSF halo.
      • Collect at least 6 TNs for each star
      • Offset out of the FoV to acquire SKY frames via the WFS control GUI with the same # of repetitions as in the TN
  • LUCI-SOUL ESM tasks - 1hr - only if no other full AO
    • Hoag's Object - on-axis AO Ref star (galaxy core as AO Ref star R=13.44 predicted WFSmag =14.7)
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • M92 or M67 at multiple PAs - separate sub-directory for these - low priority
  • if cannot switch to LUCI due to M3 SX issue
    • mid-z ULIRG FF0937+3857 - 2hrs MODS (Rothberg) - mono is OK - can collect more data at later date to improve S/N
    • LBC Solar System Objects - 1-2hrs (Veillet)
    • Local Mergers w/MODS as fill-in
  • Poor Weather Backup
    • ESM data collection (Bino-LUCIs 1hr per target)
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
      • M67 visible from handover until 03hr
      • M92 visible from 2am through twilight
      • Hoag's Object On-axis AO Ref - can run as test in poor conditions - 1 hr
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • Local Mergers (Bino-MODS or Bino-LUCI 90min or longer per target based on conditions - mono is OK too) - Rothberg
      • Any Conditions (yes even 3-4" seeing is OK)
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
      • OT file: COBandhead_Mergers_V8.xml
      • MODS
        • NGC 6598 18 08 56 69 04 04 GS R=11.8
      • LUCI
        • Arp 156 10 42 38 77 29 41 CWL=2.30 GS R=14.1 (can only observe between 22hr-02hr to get telluric, if telluric obtained before 02hr then can continue past)
        • NGC 4004 11 58 05 27 52 44 CWL=2.27 GS R=14.2
        • UGC 9829 15 23 01 -01 20 50 CWL=2.30 GS R=13.0
    • H-band Library QSO Test
      • Any conditions
      • Bino or Mono
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/H-band_Library/
      • OT file: H-band_Library_V19.xml
      • PKS 1302-102 - ~45 min with overheads
        • Up all 2nd half
    • Standard Stars (only in photometric conditions - seeing is irrelevant)
      • MODS primary spectrophotometric standard in all modes
        • The table of MODS primarty spectrophotometric standards is here
        • Scripts for acq and observation (all modes) are in ~modseng/modsScripts/modsSpecPhot
      • MODS imaging standards (this is better done later - the fields are only just rising at the end of the night in March)
      • LBC imaging standards
      • LUCI Standards
        • Scripts: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/LUCI_Phot_ZP
        • OT file: LUCI_Phot_ZP_OT_V2.xml
        • Best to pick one star and sample it over different airmasses
        • Otherwise, make sure to sample airmasses over delta-airmass > 0.33


Night Log

Link to Google Sheets here

-- BarryRothberg - 02 Mar 2021
Topic revision: r4 - 04 Mar 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback